# Module-conditioned distribution of quantum circuits

### v2

@@ 3.1 BIP formulation of MS-GC for $k\geq 4$
which denotes the latest time a unary gate is applied to $q_{i}$ up to time $t$ (or 00 if none exists). We define the following sets of migrations:
[…]
$(q_{j},\pi(q_{i}),f(j,t))\},$
[…]
$\mid p\in P\setminus\{\pi(q_{i}),\pi(q_{j})\}\}.$
@@
$M^{(1)}$ $\equiv\bigcup_{g\in B}M^{(1,g)},\quad M^{(2)}\equiv\bigcup_{g\in B}M^{(2,g)},$
@@
In order to describe a BIP formulation of MS-GC, we now view these sets as lists with indices as subscripts such that for any set $S$ we can employ a slight abuse of notation:
@@
Any ordering is acceptable as long as one remains consistent with it.
We begin by describing a BIP formulation of MS-GC for $k\geq 4$. A more efficient formulation for $k=3$ is described in Section 3.2. Consider the binary matrix
@@
Each row of $A$ contains $2+(k-2)=k$ non-zero elements. Let $x\in\{0,1\}^{|\tilde{M}|}$ represent the set of all selected migrations and pairs of migrations, which is a subset of $\tilde{M}$. To ensure that $x$ is a valid representation, we impose the constraint that if a pair of migration is selected, then each migration in the pair must also be selected. For each pair of migrations $\textbf{m}\in M^{(2)}$, we add a constraint to the vector $x$ as $e^{(\textbf{m})\top}x\geq 0$, where
@@
and $e_{i}$ denotes the $i$-th standard unit vector in $|\tilde{M}|$ dimensions.
@@
and $A$ is defined as in (1) (see Algorithm 1). We explain the validity of this formulation through the following lemma.
Proposition 1.
A solution of (3.1) is an optimal solution of MS-GC.
Proof.
We show that it suffices to relax the constraint
to
Since $x$ is a binary vector, (4) is equivalent to
and these two constraints can be merged into a single constraint as
reducing the number of constraints by half. The equivalence of (3.1) and (6) is verified in Table 1.
Algorithm 1 BIP for MS-GC, $k\geq 4$
2:  Output: Vector of selected migrations $x$
3:  Initialize:
4:   $v^{\top}\leftarrow\left[\bm{1}_{|M|}^{\top}\bm{0}_{|M^{(2)}|}^{\top}\right]$
5:   $\tilde{A}\in\mathbb{B}^{\left(|B|+|M^{(2)}|\right)\times|\tilde{M}|}\leftarrow 0$
6:   $b^{\top}\leftarrow\left[\bm{1}_{|B|}^{\top}\bm{0}_{|M^{(2)}|}^{\top}\right]$
7:  for $j=1$ to $|\tilde{M}|$ do
8:     for $i=1$ to $|B|$ do
9:        if $\tilde{M}_{j}\in\tilde{M}^{(B_{i})}$ then
10:           $\tilde{A}_{ij}\leftarrow 1$
11:        end if
12:     end for
13:     for $i=1$ to $|M^{(2)}|$ do
14:        if $\tilde{M}_{j}\in M^{(2)}_{i}$ then
15:           $\tilde{A}_{(|B|+i)j}\leftarrow 1$
16:        else if $\tilde{M}_{j}=M^{(2)}_{i}$ then
17:           $\tilde{A}_{(|B|+i)j}\leftarrow-2$
18:        end if
19:     end for
20:  end for
21:  Solve BIP for $x\in\mathbb{B}^{|\tilde{M}|}$:
22:   Minimize $v^{\top}x$ subject to $\tilde{A}x\geq b$.
23:  return $x$
Table 1: Truth table for (3.1) & (6).
$x_{\tilde{M}.\text{idx}(\textbf{m}_{1})}$ $x_{\tilde{M}.\text{idx}(\textbf{m}_{2})}$ $x_{\tilde{M}.\text{idx}(\textbf{m})}$ (3.1) (6)
@@
The only combination that violates (3) while making (6) true is
@@
However, it is easy to see that this does not disturb the formulation. Let $\hat{x}$ be an optimal solution of (3.1) and assume that
@@
i.e., $\hat{x}^{(\textbf{m})}$ is also an optimal solution of (3.1), which satisfies (3). ∎
@@
3:  Initialize:
4:   $v^{\top}\leftarrow\bm{1}_{|M^{(1)}|}^{\top}$
5:   $A\in\mathbb{B}^{|B|\times|M^{(1)}|}\leftarrow 0$
6:   $b\leftarrow 2\cdot\bm{1}_{|B|}$
7:  for $i=1$ to $|B|$ do
8:     for $j=1$ to $|M^{(1)}|$ do
9:        if $M_{j}^{(1)}\in M^{(1,B_{i})}$ then
10:           $A_{ij}\leftarrow 2$
11:        else if $M_{j}^{(1)}\in M^{(2,B_{i})}_{1}$ then
12:           $A_{ij}\leftarrow 1$
13:        end if
14:     end for
15:  end for
16:  Solve BIP for $x\in\mathbb{B}^{|M^{(1)}|}$:
17:   Minimize $v^{\top}x$ subject to $Ax\geq b$.
18:  return $x$
@@ 3.2 BIP formulation of MS-GC for $k=3$
Figure 3: Gray boxes represent module indices. (a) Removing the red migration leaves the red gate uncovered, which is covered by the blue migration. (b) Removing the red migration leaves the red gates uncovered, which is not covered by any single migration.
@@
where $g=(\{q_{i},q_{j}\},t)$ and $p\in P\setminus\{\pi(q_{i}),\pi(q_{j})\}$ is unique. Before we describe the objective function, we remark that the optimization is respect to $x\in\mathbb{B}^{|M^{(1)}|}$ instead of $x\in\mathbb{B}^{|M|}$. We justify this reduction of variables through the following lemma.
@@
We prove by contradiction. Let $P=\{p_{1},p_{2},p_{3}\}$ and $\pi(q_{i})=p_{1}$. Suppose $m=(q_{i},p_{2},t)\notin M^{(1)}$ for some $m\in M$ is strictly necessary. By the definition of $M^{(1)}$, there is no gate ($(\{q_{i},q_{j}\},t^{\prime})$ such that $\pi(q_{j})=p_{2}$ and $f(i,t^{\prime})=t$. This means that the non-local binary gates that $m$ can cover (by itself or in a pair with another migration) must be in the form of ($(\{q_{i},q_{j}\},t^{\prime})$, where $\pi(q_{j})=p_{3}$ and $f(i,t^{\prime})=t$ (and in fact, $m$ cannot cover one of these gates by itself because it migrates $q_{i}$ in $p_{1}$ to $p_{2}$ while the gate acts on a qubit in $p_{3}$; it must be paired with a migration that sends $q_{j}$ to $p_{2}$). Since we assumed that $m$ is strictly necessary, there must exist at least one such gate. Meanwhile, notice that those gates are all covered by the single migration $m^{\prime}=(q_{i},p_{3},t)\in M^{(1)}$. Therefore, replacing $m$ with $m^{\prime}$ does not increase the number of selected migrations without introducing any uncovered gates, which implies that any migration $m\notin M^{(1)}$ is not strictly necessary; each of them can be replaced by some $m^{\prime}\in M^{(1)}$. ∎
Fig. 3 provides an example circuit to visualize Lemma 1, as well as a circuit with $k=4$ to which the lemma does not apply.
@@
where $A$ is defined as in (7). We explain the validity of this formulation through the following lemma.
Proposition 2.
A solution of (3.2) is an optimal solution of MS-GC for $k=3$.
@@
which does not cover $g$. In general, it is impossible for four positive ‘rewards’ $r_{1}$, $r_{2}$, $r_{3}$, and $r_{4}$ to satisfy the conditions where $r_{1}+r_{2}$ and $r_{3}+r_{4}$ are greater than or equal to a threshold th, while $r_{1}+r_{3}$, $r_{1}+r_{4}$, $r_{2}+r_{3}$, and $r_{2}+r_{4}$ are all less than th.
3.3 Examples and analysis
Fig. 4 shows an example circuit distributed by the hypergraph partitioning formulation introduced in Section 2.4 and the BIP formulation described in this section. Even for these small circuits, the hypergraph partitioner begins to find suboptimal solutions, and the BIP post-processing helps reduce the ebit cost.
@@ 3.3 Examples and analysis
for (3.2). Also, the number of constraints is
@@
for (3.1) and $|B|$ for (3.2).
Remark 1.
The BIP formultations (3.1) and (3.2) is naturally generalized to heterogeneous settings, where the communication cost between each pair of modules is not constant. We simply replace $\left[\bm{1}_{|M|}^{\top}\bm{0}_{|M^{(2)}|}^{\top}\right]$ and $\bm{1}_{|M^{(1)}|}^{\top}$ with the vectors that represent the cost function.
Figure 4: (a) Distribution found by a hypergraph partitioner and (b) module-conditioned BIP. Numbers denote qubit indices and wires of the same color are assigned to the same module.
@@ 3.4 Partitioning QFT
In Section 3, we have discussed the suboptimality of Ref. [2] for MS-GC given a module allocation function $\pi$. Meanwhile, we observe that it also returns suboptimal module allocation functions (recall that the hypergraph partitioning formulation finds $\pi$ and selects migrations at once).
Assuming an arbitrary module allocation and MS-GC, finding an optimal $\pi$ for a given circuit is generally a hard problem. Here, we explain a very specific case that can be solved easily. That is, assuming a balanced module allocation and MS-HC, we can find an optimal $\pi$ for a QFT circuit.
@@
Figure 5: $5$-qubit QFT circuit (SWAP gates omitted).
Note that we can omit the final SWAP gates and interpret the qubits in reverse order [27, 30]. In the subsequent diagrams, we replace each $H$/$CP(2\pi/2^{k})$ gate with a general unary/$CP$ gate symbol.
It turns out that we can derive an optimal solution for any $(k,m)$-balanced distribution of QFT, assuming MS-HC.
@@
{($\left\{\left(\{q,q^{\prime}\},t_{\{q,q^{\prime}\}}\right)\mid\pi(q^{\prime})=p\right\},$
@@
[…]
$=\binom{n}{2}-k\binom{m}{2}$
@@
[…]
$\,\lceil j/m\rceil=\kappa\Big{\}}$ (11)
and the circuit do distribute is $C=U\cup B$. But each ($\left(\{q_{i},q_{j}\},t_{\{q_{i},q_{j}\}}\right)$ in (3.4) is covered by
@@
covers all gates in $B$ with ebit cost $k(k-1)m/2$, which achieves its lower bound (9). ∎
Figure 6: (a) A $6$-qubit QFT circuit. (b–d) Circuit $C$, MS-HC, and MS-GC for module allocation ‘$112233$’. (e–g) Circuit $C$, MS-HC, and MS-GC for module allocation ‘$122331$’. Both (c) and (f) involve $6$ migrations, which is optimal for MS-HC. However, (d) involves $4$ migrations while (g) involves $5$ migrations for MS-GC.
@@
Figure 7: Histograms of ebit cost for QFT circuits. Height indicates the number of allocations.
Note that $\pi^{*}$ in Lemma 2 is not the only optimal module allocation for MS-HC (see Fig. 6). However, for $(k,m)=(3,2)$, it is very likely that $\pi^{*}$ is the only optimal module allocation for MS-GC (see Table 2). We also ran exhaustive experiments to find the optimal allocations among all balanced allocations for $(k,m)=(4,2)$ and $(k,m)=(3,3)$ (see Fig. 7). In both cases, the canonical partitions $\pi^{*}$ were optimal for MS-GC (assuming that the BIP solver found optimal solutions for these small examples). That is, distributing a QFT circuit by solving the BIP associated with $\pi^{*}$ is expected to yield the lowest ebit cost we can hope for. Indeed, we observe that this heuristic proves effective (see Section 4.2.1).
@@
One might be tempted to apply this heuristic to general circuits, i.e., finding an optimal $\pi$ for MS-HC and solving the associated BIP for MS-GC. Unfortunately, this approach appears impractical at present because in general, it is unlikely that we can find an optimal $\pi$ even for MS-HC. Also, there is no guarantee that an optimal $\pi$ for MS-HC is also optimal for MS-GC.
4 Experiments
The same algorithm and configuration described in Ref. [1] were used for hypergraph partitioning. For the BIP solver, we used Gurobi 11.0.0 with a Named-User Academic License [18]. It is based on the branch-and-bound paradigm for discrete optimization problems [9, 22]. We outline the types of circuits considered in our experiments and show the results. In all the following examples, every single run of the BIP solver resulted in an ebit cost less than or equal to the non-post-processed case, demonstrating stability and general applicability of our approach.

### v3

@@ 3.1 BIP formulation of MS-GC for $k\geq 4$
which denotes the latest time a unary gate is applied to $q_{i}$ up to time $t$ (or 0 if none exists). We define the following sets of migrations:
$M^{(1,(\{q_{i},q_{j}\},t))}$ $\equiv\{(q_{i},\pi(q_{j}),f(i,t)),(q_{j},\pi(q_{i}),f(j,t))\},$
$M^{(2,(\{q_{i},q_{j}\},t))}$ $\equiv\big\{\{(q_{i},p,f(i,t)),(q_{j},p,f(j,t))\}\mid p\in P\setminus\{\pi(q_{i}),\pi(q_{j})\}\big\}.$ (1)
@@
$M^{(1)}$ $\equiv\bigcup_{g\in B}M^{(1,g)},\quad M^{(2)}\equiv\bigcup_{g\in B}M^{(2,g)},$ (2)
@@
To express the BIP compactly, we fix an arbitrary but consistent ordering of each set and treat it as an indexed list. This allows us to introduce decision variables and constraints using subscripts without repeatedly naming individual migrations. Specifically, for any set $S$ we employ the notation
@@
Algorithm 1 BIP for MS-GC, $k\geq 4$
2:Output: Vector of selected migrations $x$
3: Construct $M$, $M^{(2)}$, and $\tilde{M}$ using Definition 1 and Eqs. (3.1) and (3.1).
4:[Define objective and constraints]
5: Initialize:
6:  $v^{\top}\leftarrow\left[\bm{1}_{|M|}^{\top}\bm{0}_{|M^{(2)}|}^{\top}\right]$
7:  $\tilde{A}\in\mathbb{B}^{\left(|B|+|M^{(2)}|\right)\times|\tilde{M}|}\leftarrow 0$
8:  $b^{\top}\leftarrow\left[\bm{1}_{|B|}^{\top}\bm{0}_{|M^{(2)}|}^{\top}\right]$
9:[Iterate over migrations and fill coverage/consistency constraints]
[…]
11:  for $i=1$ to $|B|$ do
12:   if $\tilde{M}_{j}\in\tilde{M}^{(B_{i})}$ then
13:    $\tilde{A}_{ij}\leftarrow 1$
14:   end if
15:  end for
16:  for $i=1$ to $|M^{(2)}|$ do
17:   if $\tilde{M}_{j}\in M^{(2)}_{i}$ then
18:    $\tilde{A}_{(|B|+i)j}\leftarrow 1$
19:   else if $\tilde{M}_{j}=M^{(2)}_{i}$ then
20:    $\tilde{A}_{(|B|+i)j}\leftarrow-2$
21:   end if
22:  end for
23:end for
24:[Use an off-the-shelf solver]
25: Solve the following BIP for $x\in\mathbb{B}^{|\tilde{M}|}$:
26:   Minimize $v^{\top}x$ subject to $\tilde{A}x\geq b$.
27:return $x$
Consider the binary matrix
@@
Each row of $A$ contains 2+(k2)=k2+(k-2)=k non-zero elements. Let $x\in\{0,1\}^{|\tilde{M}|}$ represent the set of all selected migrations and pairs of migrations, which is a subset of $\tilde{M}$. To ensure that $x$ is a valid representation, we impose the constraint that if a pair of migrations is selected, then each migration in the pair must also be selected. For each pair of migrations $\textbf{m}\in M^{(2)}$, we add a constraint to the vector $x$ as $e^{(\textbf{m})\top}x\geq 0$, where
@@
[…]
@@
and $A$ is defined as in Eq. (3) (see Algorithm 1). Algorithm 1 obtains the minimizer in Step 26 by passing the BIP instance to a standard mixed-integer solver (we use Gurobi [gurobi]), which searches over $x\in\mathbb{B}^{|\tilde{M}|}$. Therefore, the returned $x$ specifies an MS-GC solution for the given $\pi$, and its objective value equals the ebit cost achievable with that solution.
To summarize, we enforce that every gate is covered by selecting at least one admissible covering option (either a home coverage or a joint coverage), while guaranteeing logical consistency by preventing the selection of a joint coverage unless its constituent migrations are also selected. The objective in Eq. (3.1) then counts the number of selected single migrations (i.e., the ebit cost), so minimizing it yields the minimum-cost feasible cover under general coverage for the fixed allocation $\pi$. Formally, we explain the validity of this formulation through the following observation.
Table 1: Truth table for Eqs. (3.1) and (8).
$x_{\tilde{M}.\text{idx}(\textbf{m}_{1})}$ $x_{\tilde{M}.\text{idx}(\textbf{m}_{2})}$ $x_{\tilde{M}.\text{idx}(\textbf{m})}$ Eq. (3.1) Eq. (8)
@@
Observation 1.
A solution of Eq. (3.1) is an optimal solution of MS-GC.
Proof.
We show that it suffices to relax the constraint
to
Since $x$ is a binary vector, Eq. (6) is equivalent to
and these two constraints can be merged into a single constraint as
reducing the number of constraints by half. The equivalence of Eqs. (3.1) and (8) is verified in Table 1.
The only combination that violates Eq. (5) while making Eq. (8) true is
@@
However, it is easy to see that this does not disturb the formulation. Let $\hat{x}$ be an optimal solution of Eq. (3.1) and assume that
@@
i.e., $\hat{x}^{(\textbf{m})}$ is also an optimal solution of Eq. (3.1), which satisfies Eq. (5). ∎
@@
3: Construct $M^{(1,g)}$, $M^{(2,g)}$, and $M^{(1)}$ using Eqs. (3.1) and (3.1).
4:[Define objective and constraints]
5: Initialize:
6:  $v^{\top}\leftarrow\bm{1}_{|M^{(1)}|}^{\top}$
7:  $A\in\mathbb{B}^{|B|\times|M^{(1)}|}\leftarrow 0$
8:  $b\leftarrow 2\cdot\bm{1}_{|B|}$
9:[Iterate over migrations and fill coverage constraints]
10:for $i=1$ to $|B|$ do
11:  for $j=1$ to $|M^{(1)}|$ do
12:   if $M_{j}^{(1)}\in M^{(1,B_{i})}$ then
13:    $A_{ij}\leftarrow 2$
14:   else if $M_{j}^{(1)}\in M^{(2,B_{i})}_{1}$ then
15:    $A_{ij}\leftarrow 1$
16:   end if
17:  end for
18:end for
19:[Use an off-the-shelf solver]
20: Solve the following BIP for $x\in\mathbb{B}^{|M^{(1)}|}$:
21:  Minimize $v^{\top}x$ subject to $Ax\geq b$.
22:return $x$
@@ 3.2 BIP formulation of MS-GC for $k=3$
where $g=(\{q_{i},q_{j}\},t)$ and $p\in P\setminus\{\pi(q_{i}),\pi(q_{j})\}$ is unique. Before we describe the objective function, we remark that the optimization is with respect to $x\in\mathbb{B}^{|M^{(1)}|}$ instead of $x\in\mathbb{B}^{|M|}$. We justify this reduction of variables through the following lemma.
@@
We prove by contradiction. Let $P=\{p_{1},p_{2},p_{3}\}$ and $\pi(q_{i})=p_{1}$. Suppose $m=(q_{i},p_{2},t)\notin M^{(1)}$ for some $m\in M$ is strictly necessary. By the definition of $M^{(1)}$, there is no gate $(\{q_{i},q_{j}\},t^{\prime})$ such that $\pi(q_{j})=p_{2}$ and $f(i,t^{\prime})=t$. This means that the non-local binary gates that $m$ can cover (by itself or in a pair with another migration) must be in the form of $(\{q_{i},q_{j}\},t^{\prime})$, where $\pi(q_{j})=p_{3}$ and $f(i,t^{\prime})=t$ (and in fact, $m$ cannot cover one of these gates by itself because it migrates $q_{i}$ in $p_{1}$ to $p_{2}$ while the gate acts on a qubit in $p_{3}$; it must be paired with a migration that sends $q_{j}$ to $p_{2}$). Since we assumed that $m$ is strictly necessary, there must exist at least one such gate. Meanwhile, notice that those gates are all covered by the single migration $m^{\prime}=(q_{i},p_{3},t)\in M^{(1)}$. Therefore, replacing $m$ with $m^{\prime}$ does not increase the number of selected migrations without introducing any uncovered gates, which implies that any migration $m\notin M^{(1)}$ is not strictly necessary; each of them can be replaced by some $m^{\prime}\in M^{(1)}$. ∎
Figure 4: Illustration of why $k=3$ admits a reduced formulation (Lemma 1) while $k\geq 4$ does not. Gray boxes show module indices (each module contains a single qubit), and arrows denote migrations. (a) $k=3$: Removing the red migration makes the red gate uncovered, but it can still be covered by the single blue migration because the blue migration (which is an element of $M^{(1)}$) moves the same qubit to the other endpoint’s home module. This reflects that migrations outside $M^{(1)}$ can be replaced without increasing cost. (b) $k=4$: The circuit includes an additional set of gates, and removing the red migration leaves the red gates uncovered in a way that cannot be repaired by any single migration. This contrast explains why we can restrict migrations to $M^{(1)}$ for $k=3$, but not for $k\geq 4$.
Fig. 4 provides an example circuit to visualize Lemma 1, as well as a circuit with $k=4$ to which the lemma does not apply. The contrast stems from the fact that when $k=3$ there is only one “third” module besides the two endpoint-modules of any migration, whereas for $k\geq 4$ there are multiple possible third modules.
@@
where $A$ is defined as in Eq. (9). We explain the validity of this formulation through the following observation.
Observation 2.
A solution of Eq. (3.2) is an optimal solution of MS-GC for $k=3$.
@@
which does not cover $g$. In general, it is impossible for four positive “rewards” $r_{1}$, $r_{2}$, $r_{3}$, and $r_{4}$ to satisfy the conditions where $r_{1}+r_{2}$ and $r_{3}+r_{4}$ are greater than or equal to a threshold th, while $r_{1}+r_{3}$, $r_{1}+r_{4}$, $r_{2}+r_{3}$, and $r_{2}+r_{4}$ are all less than th.
Remark 1.
The BIP formulations in Eqs. (3.1) and (3.2) are naturally generalized to heterogeneous settings, where the communication cost between each pair of modules is not constant. We simply replace $\left[\bm{1}_{|M|}^{\top}\bm{0}_{|M^{(2)}|}^{\top}\right]$ and $\bm{1}_{|M^{(1)}|}^{\top}$ with the vectors that represent the cost function.
3.3 Illustrative example
Figure 5: (a) Distribution found by a hypergraph partitioner and (b) BIP post-processing. Numbers denote qubit indices and wires of the same color are assigned to the same module. Arrows indicate migrations. The ebit cost drops from 44 in (a) to 33 in (b). The fact that the hypergraph partitioner is suboptimal even on this tiny instance motivates our BIP post-processing step.
Selecting the module allocation together with the corresponding migrations and ebit distribution is combinatorially difficult. Even for small circuits, the hypergraph partitioner can return suboptimal distributions, and our BIP post-processing can reduce the ebit cost. For example, Fig. 5 compares the distribution produced by the hypergraph partitioning approach in Section 2.4 with the distribution after BIP post-processing: the ebit cost drops from $4$ to $3$.
3.4 Problem size of the BIP formulations
@@
for Eq. (3.2). Also, the number of constraints is
@@
for Eq. (3.1) and $|B|$ for Eq. (3.2).
@@ 3.5 Partitioning QFT
In the previous sections, we focused on the migration-selection layer. That is, for a given module allocation $\pi$, we showed that MS-GC can be formulated as a BIP. We also illustrated that the distribution produced by hypergraph partitioning [andres2019automated] can be suboptimal.
In this section, we turn to the module allocation layer itself and discuss a structured case where $\pi$ can be chosen in a principled way. We focus on QFT because it is a core subroutine used in many quantum algorithms, so improving its distributed implementation is valuable in its own right. Moreover, assuming a balanced distribution, QFT exhibits additional structure that makes identifying an optimal $\pi$ tractable in the (restricted) MS-HC setting.
Figure 6: A monolithic 66-qubit QFT circuit (SWAP gates omitted). After fixing a module allocation, we remove gates that are local to each module and optimize only the remaining non-local operations.
[…]
@@
It turns out that we can derive an optimal solution for any $(k,m)$-balanced distribution of QFT, assuming MS-HC. The resulting ebit cost coincides with that derived independently by [kaur2025optimized]. However, we strengthen their result by establishing that this cost attains a matching lower bound, thereby proving the optimality of the scheme in the MS-HC setting. Furthermore, the module allocation introduced in the proof will also be reused in the MS-GC setting.
Remark 2.
While we omit the final SWAP gates in our QFT distribution, [kaur2025optimized] achieves the same ebit cost even when these operations are explicitly included. They accomplish this by employing qubit routing, but such techniques fall outside the scope of our specific model (i.e., MS-GC under fixed $\pi$). More importantly, as noted in [neumann2020imperfect], these final SWAP gates can be omitted without introducing overhead in downstream processing. This is because the SWAP gates merely permute the logical qubits, and their effect can be compensated for through simple qubit reindexing.
In QFT, each qubit has exactly one unary gate and then interacts via $CP$ with every other qubit, so under MS-HC the natural migration time is immediately after that unary gate. In a $(k,m)$-balanced allocation, migrating a qubit into a module can cover at most $m$ gates between that qubit and the $m$ qubits in the destination module, giving a tight lower bound on the number of migrations. Formally, we have the following lemma.
@@
$\left\{\left(\{q,q^{\prime}\},t_{\{q,q^{\prime}\}}\right)\mid\pi(q^{\prime})=p\right\},$
@@
# binary# local $\#\text{ binary}$-$\text${ local binary} $=\binom{n}{2}-k\binom{m}{2}$
@@
$B=\bigcup_{i=1}^{(k-1)m}\bigcup_{\kappa=\lceil i/m\rceil+1}^{k}\Big\{\left(\{q_{i},q_{j}\},t_{\{q_{i},q_{j}\}}\right)\mid\lceil j/m\rceil=\kappa\Big\}$ (13)
and the circuit to distribute is $C=U\cup B$. But each $\left(\{q_{i},q_{j}\},t_{\{q_{i},q_{j}\}}\right)$ in Eq. (13) is covered by
@@
covers all gates in $B$ with ebit cost k(k1)m/2k(k-1)m/2, which achieves its lower bound Eq. (11). ∎
Figure 7: (a) A 66-qubit QFT circuit. (b–d) Circuit $C$, MS-HC, and MS-GC for module allocation ‘$112233$’. (e–g) Circuit $C$, MS-HC, and MS-GC for module allocation ‘$122331$’. Both (c) and (f) involve 66 migrations, which is optimal for MS-HC. However, (d) involves 44 migrations while (g) involves 55 migrations for MS-GC.
@@
Figure 8: Histograms of MS-GC ebit cost for $(k,m)$-balanced distributions of QFT, computed by enumerating all balanced module allocations for the indicated $(k,m)$ and solving the associated BIP. The bar including the canonical partition $\pi^{*}$ is highlighted in orange. In both cases, $\pi^{*}$ attains the minimum MS-GC ebit cost (after BIP) among all balanced allocations, supporting our use of $\pi^{*}$ as the default module allocation for MS-GC.
Note that $\pi^{*}$ in Lemma 2 is not the only optimal module allocation for MS-HC (see Fig. 7). However, for $(k,m)=(3,2)$, it is highly likely that $\pi^{*}$ is the only optimal module allocation for MS-GC (see Table 2). We also ran exhaustive experiments to find the optimal allocations among all balanced allocations for $(k,m)=(4,2)$ and $(k,m)=(3,3)$ (see Fig. 8). In both cases, the canonical partitions $\pi^{*}$ were optimal for MS-GC (assuming that the BIP solver found optimal solutions for these small examples). That is, distributing a QFT circuit by solving the BIP associated with $\pi^{*}$ is expected to yield the lowest ebit cost we can hope for. Indeed, we observe that this heuristic proves effective (see Section 4.2.1).
@@
One might be tempted to extend this idea beyond QFT by (i) choosing $\pi$ to be optimal for MS-HC and (ii) applying our BIP post-processing to solve MS-GC under that fixed $\pi$. In general circuits, however, Step (i) is itself a nontrivial combinatorial problem. Unlike QFT, we do not currently have a generic way to certify an optimal module allocation $\pi$ even under MS-HC, and moreover an allocation that is optimal for MS-HC need not be optimal for MS-GC. That said, there may be additional circuit families with exploitable structure where an allocation strategy can be characterized analytically or with lightweight combinatorial methods. Natural candidates include state-preparation circuits for highly symmetric target states (e.g., GHZ- and W-type constructions), as well as phase-estimation-type circuits, which contain (inverse) QFT as a subroutine. Identifying such structured classes and deriving allocation rules or provable bounds for them (and then combining these with our BIP post-processing) is an interesting direction for future work.
4 Experimental results
Experiment Ebit cost reduction (%)
Max. Min.
CZCZ fraction-10% 1.05 0.00
CZCZ fraction-30% 1.56 0.00
CZCZ fraction-50% 3.33 0.63
CZCZ fraction-70% 6.77 0.48
CZCZ fraction-90% 19.6 2.99
Quantum Volume 0.86 0.00
QFT 23.1 0.00
DraperQFTAdder 30.4 0.00
RGQFTMultiplier 95.8 0.00
AND 84.4 0.00
InnerProduct 85.7 0.00
Table 3: Summary of the ebit-cost savings achieved by BIP post-processing on top of the hypergraph-partitioning baseline (HP). For each benchmark family, we report the maximum and minimum percentage reduction in total ebit cost over all tested configurations.
The same algorithm and configuration described in Ref. [andres2024distributing] were used for hypergraph partitioning. For the BIP solver, we used Gurobi 11.0.0 with a Named-User Academic License [gurobi]. It is based on the branch-and-bound paradigm for discrete optimization problems [clausen1999branch, land2009automatic]. Table 3 outlines the types of circuits considered in our experiments and summarizes the results by reporting the maximum and minimum ebit-cost reduction (in %) across all tested configurations for each family. Full results for every parameter setting are deferred to Appendix D. Notably, the minimum reduction is nonnegative in all cases, i.e., the BIP post-processing never produces a worse ebit cost than the HP baseline in any run.
Remark 4.
Our BIP post-processing is performed during offline distribution and therefore adds classical runtime compared to a baseline solver. However, since the goal is to reduce execution-time ebit consumption, this overhead is often amortized when a distributed circuit is run many times. In such cases, the one-time post-processing overhead may be outweighed by the repeated savings in ebit cost during execution. Moreover, our method reduces MS-GC to a BIP and then delegates optimization to a standard solver (Gurobi). In the current implementation, the largest instances solved in our environment are around $n=50$, $k=8$. The current scalability ceiling is primarily driven by solver capability on these instances. We plan to extend this regime by exploiting circuit-level local regularities to eliminate redundant variables and constraints and thereby accelerate the optimization, while leaving improvements to general-purpose BIP technology outside the scope of this manuscript.
In what follows, we describe the experimental circuit families and their structures in detail. These benchmarks are largely divided into two categories: families of random circuits and families of structured arithmetic circuits. For each family we provide a representative circuit diagram.